Your best friend just won the Florida lottery. She has the choice of $15.900,000 today or an annuity with payments of $1.050.000 for 20 years, with the first payment coming one year from today. What rate of return is built into the annuity? Disregard any tax consequences. Your answer should be between 2.01 and 14.74, rounded to 2 decimal places, with no special characters. Question 12 5 pts lim sold a car and accepted a note promising cash flows of $1 000 at the end of Year 1. and $2 000 at the end of Vears 2 and 4 as his payment. What was the effective price he received for the car assuming an interest rate of 7.10%? Your answer should be between 4 /45,oo and 6.525.00, rounded to 2 decimal places with no spedial charachers, tetie
Answer #1

Question 1 - Answer is 2.81%

Solving rate of return can only be done using Excel or Financial calculation. For PV of annuity formula, you cannot solve for rate of return manually. Here is the formula:

P Periodic Payment r-rate per period n-number of periods

Try, putting values in this and solving for r to understand why we can't solve it manually. Here is the Excel snapshot (with calculator inputs as in bracket):

Question 2: Answer is $5,825.45

We need to calculate the PV of all the cash flows that would be made in order to calculate the effective price of car.

PV = FV/(1 + r)n

PV = \frac{1000}{(1 + 0.071)^1} + \frac{2000}{(1 + 0.071)^2} + \frac{2000}{(1 + 0.071)^3} + \frac{2000}{(1 + 0.071)^4}

PV = 933.71 + 1743.62 + 1628.03 + 1520.10

PV = $5,825.45